# ThinkRec: Thinking-based recommendation via LLM

### v3

@@ Appendix B Prompt Templates
B.1 Prompt for Summarizing Metadata
@@
B.2 Prompt for Recommendation
Recommendation
#Question: A user has given ratings to the following books: <HisItemList>. Additionally, we have information about the user’s preferences encoded in the feature <UserID>. Based on the descriptions and the user’s enjoyment of each book in the historical sequence, construct a persona of the user’s preferences and reevaluate whether the user would enjoy the book titled <TargetItemTitle> with the feature <TargetItemID>. Please begin your analysis with "Yes" or "No".
#Answer:
B.3 Prompt for Reasoning Data Synthesis
@@
A user has given high ratings to the following books: <HisItemList>. Using all available information, make a prediction about whether the user would enjoy the book titled <TargetItemTitle>?
@@
The correct response is <answer>. Reflect on multiple aspects based on historical information and explain the reason for the oversight based on the previous analysis. Reanalyze to make a prediction about whether the user would enjoy the book titled <TargetItemTitle>?
@@ Appendix C Case Study
Among existing LLM-based recommendations, CoLLM and Prompt4NR yield disorganized symbols. TALLRec frequently generates sentences with unrelated elements, such as code or hallucinated facts, failing to reflect coherent reasoning. In contrast, ThinkRec demonstrates structured, step-by-step reasoning aligned with user history and target item semantics, enabling it to produce accurate and interpretable recommendations.
@@
")) ’"’); 3.}"’, ’); }’) , } "); }; """ “‘ I think I understand the problem, but it’s a bit unclear to me. ") }; } """ ");
@@
");" # 3. The user’s response: "1"; (the number of the book to be analyzed) # - If a book has been rated, then it is included in this list. # * For each book that matches with any other book on the same list: # + Check if both books are novels and have similar genres # - In case they match these criteria, add them together as one single entry """ def predict_book_rating(book_id): for i in range(len(user_ratings)): if int(i) == book_id: return True result = [] for j in user_ratings: if len(result) > 0: flag = False for k in result[-1]: if j[book_name] == k[0][book_title]: if ’novel’ in j[’genres’] or ’fiction’ or ”: if (’romance’, ”) not in k: result.pop() break else: flag=True continue elif j.get(’genres’) != None and set(j[’genre’]) & set(k[2]): flag=False # #!/usr/bin/env python from flask import Flask, request, jsonify import json app = Flask(__name__) @app.route(’/predict-book-rating’) def predict(): data = {’user’: [’The Duchess Deal: Girl Meets Duke’; ’The V Card’;…]} print(data) if __name__ == ’__main__’: app.run(debug=True) “‘ Here you can use Python script. Please note that I am using ‘flask‘ framework which allows us to create simple web applications. Here we will make an API endpoint ‘/predict-books-ratings‘. You should send JSON object where key ‘’books’‘ contains array of book titles.
@@
"); #</#">’; ’ ’, ’; ’# ’; “‘" I have been trying to solve this problem for a long time, and I am still unable. Can you help me?
@@
. Okay, let’s see here. The user has rated several books positively and I need to figure out if they’d like the new book "If I Run (If i Run Series)". First, looking at their liked books: They enjoy thrillers with elements of suspense, crime, mystery, family dynamics, abuse themes, and some historical or biographical aspects too. For example, in "A Child Called It" there’s childhood trauma, healing, which is a strong theme. Then series like Heist and Millennium have action, romance maybe? But mostly thriller/suspense. Even though it mentions romance in those, the main focus seems on the heists and cybercrime. The next one, "Saving Parker", includes American history and children’s angles again. So historical fiction blended with personal stories might be appealing. Also, Jack Reacher novels are part of this list, so that suggests they like gritty detective work or military/authoritative figures. Michael Bennett and Alex Cross also indicate police procedural or investigative mysteries. Now the target book is an If I’m Found sequel, focusing on family, suspense/mystery, murder, love, marriage/divorce… If you run has plot twist, dialogue, compulsive reading, etc., which aligns with what users like from other books. Themes include family and suspense. Their existing likes include both family-focused (e.g., Saving Parker) along with intense plots (Heist, Stieg Larsson). The presence of murder and investigation matches Haunted by Michael Bennet. Marriage/divorced themes aren’t explicitly mentioned elsewhere, but relationships are present in others (Before We… has love story?). So overall, given the overlap in key genre tags (family, thriller, suspension), plus positive reviews on related topics, I think the answer should be Yes.
Appendix D Notation
Symbol Description
@@ Appendix D Notation
$u,t$ User and time indices
$N$ Total number of users
$T_{u}$ Number of behaviors of user $u$
=$\mathcal{E}=\{\mathbf{e}^{c}_{1:N}\}$ Representations of experts
@@
$i_{id},i_{txt},i_{t},i_{d},i_{k},i_{l}$ Item ID, textual information, title, description, keywords, yes/no label
@@
$\mathbf{emb}_{s}^{u},\mathbf{emb}_{s}^{i}$ User/item embeddings projected into language space
@@
$L$ Number of textual tokens
$pos$, $posid$ Position of answer token; token ID of “Yes”
$\tau$ Softmax temperature
@@
Table 5: Summary of Notations

### v4

@@ Appendix B Prompt Templates
B.1. Prompt for Summarizing Metadata
@@
B.2. Prompt for Reasoning Data Synthesis
@@
A user has given high ratings to the following books: ¡HisItemList¿. Using all available information, make a prediction about whether the user would enjoy the book titled ¡TargetItemTitle¿?
@@
The correct response is ¡answer¿. Reflect on multiple aspects based on historical information and explain the reason for the oversight based on the previous analysis. Reanalyze to make a prediction about whether the user would enjoy the book titled ¡TargetItemTitle¿?
@@
B.3. Prompt for Recommendation
Recommendation
#Question: A user has given ratings to the following books: ¡HisItemList¿. Additionally, we have information about the user’s preferences encoded in the feature ¡UserID¿. Based on the descriptions and the user’s enjoyment of each book in the historical sequence, construct a persona of the user’s preferences and reevaluate whether the user would enjoy the book titled ¡TargetItemTitle¿ with the feature ¡TargetItemID¿. Please begin your analysis with ”Yes” or ”No”.
#Answer:
@@ Appendix C Case Study
”)) ’”’); 3.}”’, ’); }’) , } ”); }; ””” “‘ I think I understand the problem, but it’s a bit unclear to me. ”) }; } ””” ”);
@@
”);” # 3. The user’s response: ”1”; (the number of the book to be analyzed) # - If a book has been rated, then it is included in this list. # * For each book that matches with any other book on the same list: # + Check if both books are novels and have similar genres # - In case they match these criteria, add them together as one single entry ””” def predict_book_rating(book_id): for i in range(len(user_ratings)): if int(i) == book_id: return True result = [] for j in user_ratings: if len(result) ¿ 0: flag = False for k in result[-1]: if j[book_name] == k[0][book_title]: if ’novel’ in j[’genres’] or ’fiction’ or ”: if (’romance’, ”) not in k: result.pop() break else: flag=True continue elif j.get(’genres’) != None and set(j[’genre’]) & set(k[2]): flag=False # #!/usr/bin/env python from flask import Flask, request, jsonify import json app = Flask(__name__) @app.route(’/predict-book-rating’) def predict(): data = {’user’: [’The Duchess Deal: Girl Meets Duke’; ’The V Card’;…]} print(data) if __name__ == ’__main__’: app.run(debug=True) “‘ Here you can use Python script. Please note that I am using ‘flask‘ framework which allows us to create simple web applications. Here we will make an API endpoint ‘/predict-books-ratings‘. You should send JSON object where key ‘’books’‘ contains array of book titles.
@@
”); #¡/#”¿’; ’ ’, ’; ’# ’; “‘” I have been trying to solve this problem for a long time, and I am still unable. Can you help me?
@@
. Okay, let’s see here. The user has rated several books positively and I need to figure out if they’d like the new book ”If I Run (If i Run Series)”. First, looking at their liked books: They enjoy thrillers with elements of suspense, crime, mystery, family dynamics, abuse themes, and some historical or biographical aspects too. For example, in ”A Child Called It” there’s childhood trauma, healing, which is a strong theme. Then series like Heist and Millennium have action, romance maybe? But mostly thriller/suspense. Even though it mentions romance in those, the main focus seems on the heists and cybercrime. The next one, ”Saving Parker”, includes American history and children’s angles again. So historical fiction blended with personal stories might be appealing. Also, Jack Reacher novels are part of this list, so that suggests they like gritty detective work or military/authoritative figures. Michael Bennett and Alex Cross also indicate police procedural or investigative mysteries. Now the target book is an If I’m Found sequel, focusing on family, suspense/mystery, murder, love, marriage/divorce… If you run has plot twist, dialogue, compulsive reading, etc., which aligns with what users like from other books. Themes include family and suspense. Their existing likes include both family-focused (e.g., Saving Parker) along with intense plots (Heist, Stieg Larsson). The presence of murder and investigation matches Haunted by Michael Bennet. Marriage/divorced themes aren’t explicitly mentioned elsewhere, but relationships are present in others (Before We… has love story?). So overall, given the overlap in key genre tags (family, thriller, suspension), plus positive reviews on related topics, I think the answer should be Yes.
Appendix D Implementation Details
We implement all the compared methods using PyTorch 2.5. We adopt BCE when not otherwise specified. LLM-based methods are optimized using AdamW, while other models use Adam.
Regarding traditional models, we explore the learning rate within the range of [1e-2, 1e-3, 1e-4], the embedding size within the range of [64, 128, 256], and weight decay within the range of [1e-2, 1e-3, . . . , 1e-7]. As for LLM-based methods, we set the learning rate to 1e-4, and weight decay to 1e-3 to align with CoLLM. For SASRec, we establish the maximum length of historical interaction sequences to 25. We adopt TALLRec’s practice of setting the maximum sequence length to 10 for all other methods.
Regarding other specific parameters of the baseline models, we adhere to the configurations outlined in their original papers. For the LoRA module, we follow the same configuration as CoLLM, setting r,alpha,dropout,targetmodulesr,alpha,dropout,target\ modules to 8, 16, 0.05, and “[q proj, v proj]”, respectively. We set the sample rate of reasoning data and recommendation data to 0.2 and 0.8, and the weight of each loss function as follows: $\alpha=0.1,\beta=0.9,\eta=0.9,\gamma=0.1$. We set the temperature coefficient $\tau$ to $0.1$.
Appendix E Notation
Table 7. Summary of Notations
[…]
@@ Appendix E Notation
[…]
    NN     Total number of users
[…]
    $\mathcal{E}=\{\mathbf{e}^{c}_{1:N}\}$     Representations of experts
@@
[…]
@@
[…]
@@
[…]
    pospos, posidposid     Position of answer token; token ID of “Yes”
[…]
@@
    Functions
[…]
    $\mathrm{TKZ}(\cdot)$     Tokenizer of LLM
[…]
    $proj_{\phi}(\cdot)$     Projection into language space
[…]
    $\mathrm{Length}(\cdot)$     Sequence length
[…]
    $\mathrm{BCE}(\cdot,\cdot)$     Binary cross-entropy loss
[…]
    $\mathrm{Cosim}(\cdot,\cdot)$     Cosine similarity
[…]
    Loss Functions
[…]
    $\mathcal{L}_{rec}$     Loss for the recommendation task
[…]